OpenStudy (henrietepurina):

Part A: Sam rented a boat at $225 for 2 days. If he rents the same boat for 5 days, he has to pay a total rent of $480. Write an equation in the standard form to represent the total rent (y) that Sam has to pay for renting the boat for x days. (4 points) Part B: Write the equation obtained in Part A using function notation.(2 points) Part C: Describe the steps to graph the equation obtained above on the coordinate axes. Mention the labels on the axes and the intervals. (4 points)

OpenStudy (kc_kennylau):

So we'll write it in point-slope form first, then convert it into standard form.

OpenStudy (henrietepurina):

ok so...

OpenStudy (kc_kennylau):

So find out the slope first

OpenStudy (kc_kennylau):

The slope of a function is rise over run, or y-distance/x-distance, or in this context "the increase of rent over the increase of number of days"

OpenStudy (henrietepurina):

255 over 2

OpenStudy (henrietepurina):

or 127.5 over 1?

OpenStudy (kc_kennylau):

Please be noted that I'm talking about "the increase" instead of the actual amount

OpenStudy (kc_kennylau):

From $225 for 2 days to $480 for 5 days

OpenStudy (kc_kennylau):

What's the increase of rent over the increase of number of days?

OpenStudy (henrietepurina):

255 over 3

OpenStudy (henrietepurina):

or 85 per day?

OpenStudy (kc_kennylau):

Yep

OpenStudy (henrietepurina):

So the answer for a is 85

OpenStudy (kc_kennylau):

Nope, 85 is just the slope of function

OpenStudy (henrietepurina):

so umm =85x+y

OpenStudy (kc_kennylau):

Nope, the equation given by a function with known slope \(m\) and a known point \(x_1,y_1\) is given by: \[y-y_1=m(x-x_1)\]

OpenStudy (henrietepurina):

so umm y - y = 85(x - x) i cant write the small 1

OpenStudy (kc_kennylau):

No, \(x_1,y_1\) is a point on the function

OpenStudy (kc_kennylau):

You can use (2,225) or (5,480)

OpenStudy (henrietepurina):

225 - 480 = 85(2 - 5)

OpenStudy (kc_kennylau):

The \(x\) and \(y\) are not to be substituted

OpenStudy (henrietepurina):

ohh so y = 225 = 85(x - 2)

OpenStudy (kc_kennylau):

exactly

OpenStudy (kc_kennylau):

Now convert it into standard form

OpenStudy (henrietepurina):

what was it?

OpenStudy (kc_kennylau):

Just expand everything and put everything on the left hand side of the equation

OpenStudy (henrietepurina):

how do you expand

OpenStudy (kc_kennylau):

I mean distribute

OpenStudy (henrietepurina):

how?

OpenStudy (henrietepurina):

like umm 225 = 85x - 2

OpenStudy (kc_kennylau):

y - 225 = 85(x - 2) y - 225 = 85x - 170 y = 85x - 170 + 225 y = 85x + 55 0 = 85x - y + 55 85x - y + 55 = 0

OpenStudy (henrietepurina):

i might have done this ine thing inccorectly 25 = 85x + 2

OpenStudy (henrietepurina):

YOOHOO?

OpenStudy (henrietepurina):

and btw we got this, and 2 more and they have to be done by 12:30

OpenStudy (kc_kennylau):

If you don't get the concept, useless for us to give you the answer

OpenStudy (henrietepurina):

Still, please help me

OpenStudy (kc_kennylau):

85x - y + 55 = 0

OpenStudy (henrietepurina):

you sure?

OpenStudy (henrietepurina):

i dont understand the 0 at the end

OpenStudy (henrietepurina):

hello @kc_kennylau

OpenStudy (kc_kennylau):

Which step in y - 225 = 85(x - 2) y - 225 = 85x - 170 y = 85x - 170 + 225 y = 85x + 55 0 = 85x - y + 55 85x - y + 55 = 0 did you not understand?

OpenStudy (kc_kennylau):

I can explain step by step

OpenStudy (henrietepurina):

no i just dont understand why 0 at the end!

OpenStudy (henrietepurina):

so the total cost is just 0

OpenStudy (kc_kennylau):

Nah, the rent is y

OpenStudy (henrietepurina):

oh ok

OpenStudy (henrietepurina):

Ok lets do b now

OpenStudy (kc_kennylau):

85x - y + 55 = 0 85x + 55 = y y = 85x + 55 f(x) = 85x + 55

OpenStudy (kc_kennylau):

For part C, I don't know how you'd describe it, but basically the concept is you plot the two points (2,225) and (5,480) out and then join them

OpenStudy (kc_kennylau):

I'll leave the detail to you

OpenStudy (henrietepurina):

ok thanks are you ready for 2 more?

OpenStudy (kc_kennylau):

No I have to go now

OpenStudy (henrietepurina):

ok then bye